OpenStudy (anonymous):

Wanda opened a savings account 20 years ago with a deposit of $3,150.37. The account has an interest rate of 4.1% compounded daily. How much interest has Wanda earned?

OpenStudy (anonymous):

Help please

OpenStudy (anonymous):

can any one help me

OpenStudy (tkhunny):

\(Payment * \left(1 + \dfrac{0.041}{365}\right)^{20*365} = $3,150.37\) All of this shoudl be ringing bells. Is there any piece of you you do not understand?

OpenStudy (anonymous):

thats wrong ,, these are the answers choices; $3,886.45 $7,152.59 $4,002.22 $3,279.54

OpenStudy (tkhunny):

You are right. it is wrong. I misread the given amount as the final accumulation. How did I do that? You tell me how to fix it. This is a test!

OpenStudy (tkhunny):

\(Payment \cdot\left(1+\dfrac{0.041}{365}\right)^{20*365} = Payment + Interest\) We are given the interest, so, \(Payment \cdot\left(1+\dfrac{0.041}{365}\right)^{20*365} = Payment + $3,150.37\)